OpenStudy (vicky123):

-19-7x=7x=23 . Solve for x

OpenStudy (anonymous):

Assume that's your question : -19-7x=7x+3 Right ?

OpenStudy (vicky123):

no it is -19 - 7 x = 7 x + 23

OpenStudy (jhannybean):

put the variables on one side and the constants on the other, then solve for x! :)

OpenStudy (anonymous):

Okay .First of all you should compile all like terms on one side and combine them : -19-23=7x+7x ====>(the variables on one side (Number which contain (x) ) -42=14x ==> Divide by 14 x=-3 <==== Your answer !

OpenStudy (vicky123):

thank you :)

OpenStudy (anonymous):

\[\Huge \color{MidNightBlue}{\bf{You~Are~Welcome~!}}\]

OpenStudy (vicky123):

@Eyad 80-2x=34x+20

OpenStudy (anonymous):

Hmm,Try to solve it here and i will guide you through :) Give it a shot !

OpenStudy (vicky123):

80-20=34+2x 60=36x

OpenStudy (vicky123):

@Eyad is that is?

OpenStudy (anonymous):

\[\Huge \color{Red}{Excellent~!}\] Now divide by 36 to get the value of x :)

OpenStudy (vicky123):

1 and 24/36 @Eyad

OpenStudy (anonymous):

Good job

OpenStudy (vicky123):

when you calculate it back it is wrong

OpenStudy (anonymous):

1 and 24/36 = 5/3 So : 80-20=34x+2x 60=36x 60=36 * 5/3 60=60 Cheer !